OpenStudy (anonymous):

Which of the following could be an example of a function with a range (-infinity, a] and a domain [b , infinity) where a > 0 and b > 0? A. f(x) = -sqrt(x-b) + a B. f(x) = cubert(x+b) - a C. f(x) = -cubert(x+a) - b D. f(x) = sqrt(x-a) + b *** I will make equations more readable***

OpenStudy (anonymous):

\[A. f(x) = -\sqrt{x-b} + a\]

OpenStudy (anonymous):

\[B. f(x) = \sqrt[3]{x+b} - a\]

OpenStudy (anonymous):

\[C. f(x) = -\sqrt[3]{x+a} - b\]

OpenStudy (anonymous):

\[D. f(x) = \sqrt{x-a} + b\]

OpenStudy (anonymous):

Using what I know, I think the answer would be letter A because that is the only equations (I think) that has a negative range and a positive domain. Please help!

OpenStudy (anonymous):

A would be the answer!

OpenStudy (anonymous):

Great! Thank you!!